A charge Q is uniformly distributed throughout a sphere of radius a. Taking the potential 

at infinity as zero, the potential at r =b < a is


1
Expert's answer
2021-07-12T16:23:06-0400

Gauss law

"\smallint E.A=\frac{Q}{\epsilon_0}"

r=b<a


"E.(4\pi b^2)=\frac{\rho\frac{4\pi b^2}{3}}{\epsilon_0}"

"\rho=\frac{Q}{\frac{4\pi a^3}{3}}"

"E_{in}=\frac{qb}{4\pi\epsilon_0a^3}"

b>a

"E.A=\frac{q}{\epsilon_0}"

"E_0=\frac{q}{4\pi\epsilon_0b^2}"

Potential

"V_a-V_{infinite}=\smallint_{inf}^aE.dr"

"V_a=-\smallint_{inf}^aE.dr"

"V_a=-[\smallint_{inf}^bE_0.dr+\smallint_{a}^bE_{in}.dr]"

"V_a=-[\smallint_{inf}^b\frac{kQ}{r^2}dr+\smallint_{b}^a \frac{kQr}{R^3}dr]"

"V_{in}=\frac{kQ(3a^2-b^2)}{2a^3}"

"b=a;"

"V_{in}=\frac{3Q}{8\pi\epsilon_0 a}"

"\u2206V=V_{in}(r=a)-V_{in}(r=0)"

"∆V=\frac{-Q}{8\pi\epsilon_0 a}"